i was given this example to work with...
"if 3 apples and 2 pears cost $6.00 and 7 apples and 6 pears cost $15.00, how much will one apple and one pear cost?"
given solution:
3a+2p=6.00
7a+6p=15.00
"because these two equations do not have the same 'y' value, they must be changed in order to eliminate one of the variables."
(this is where i ran into trouble)
-3(3a+2p=6.00)=-9a - 6p=-18.00

(everything i understand perfectly up until this point. i just dont how to come to the '-3' there infront of the sum. where did it come from? and how do i know whether its even suppose to be a negative or a three for that matter? please please any help would be greatly appreciated. my final graduating year depends on my completing this last book.)

Answer by Alan3354(69443)

if 3 apples and 2 pears cost $6.00 and 7 apples and 6 pears cost $15.00, how much will one apple and one pear cost?"
given solution:
3a+2p=6.00
7a+6p=15.00
"because these two equations do not have the same 'y' value, they must be changed in order to eliminate one of the variables."
(this is where i ran into trouble)
-3(3a+2p=6.00)=-9a - 6p=-18.00
-----------------
3a+2p=600
7a+6p=1500
Multiply the 1st eqn by -3
-9a - 6p = -1800
+7a + 6p = 1500
----------------- Add
-2a = -300
a = 150 cents, etc.
=========================
You were doing the right thing, just didn't finish it.